A surveyor measured the angle of elevation of a flat spire as from a point on horizontal ground. He moves 30m nearer to the flat and measures the angle of elevation as . Calculate the height of the spire to the nearest hundredth.


1
Expert's answer
2021-03-29T14:17:42-0400

tan"\theta" = "\frac{h-a}{d}"


d = 30 m

h = height of spire

"\theta" = angle of elevation

a = height of surveyor


tan"\theta" = "\frac{h-a}{30}"


30×tan"\theta" = h - a

h = (30 × tan"\theta" ) + a